# Efficient Tensor Completion Algorithms for Highly Oscillatory Operators

### v2

@@ Abstract
This paper presents low-complexity tensor completion algorithms and their efficient implementation to reconstruct highly oscillatory operators discretized as $n\times n$ matrices. The underlying tensor decomposition is based on the reshaping of the input matrix and its butterfly decomposition into an order $\mathcal{O}(\log n)$ tensor. The reshaping of the input matrix into a tensor allows for representation of the butterfly decomposition as a tensor decomposition with dense tensors. This leads to efficient utilization of the existing software infrastructure for dense and sparse tensor computations. We propose two tensor completion algorithms in the butterfly format, using alternating least squares and gradient-based optimization, as well as a novel strategy that uses low-rank matrix completion to efficiently generate an initial guess for the proposed algorithms. To demonstrate the efficiency and applicability of our proposed algorithms, we perform three numerical experiments using simulated oscillatory operators in seismic applications. In these experiments, we use $\mathcal{O}(n\log n)$ observed entries in the input matrix and demonstrate an $\mathcal{O}(n\log^{3}n)$ computational cost of the proposed algorithms, leading to a speedup of orders of magnitudes per iteration for large matrices compared to the low-rank matrix and quantized tensor-train completion. Moreover, the proposed butterfly completion algorithms, equipped with the novel initial guess generation strategy, achieve reconstruction errors that are smaller by an order of magnitude, enabling accurate recovery of the underlying structure compared to the state-of-the-art completion algorithms.
@@ 1 Introduction
Matrix or tensor completion is the task of inferring missing entries in 2D matrices or higher-order tensors from partial observations, which may be corrupted by noise. Completion problems have been widely studied for applications across data science (e.g., recommendation system [23] and computer vision [55]), imaging science (e.g., seismic inversion [54, 47], magnetic resonance imaging [90, 53], and X-ray tomography [9]), and high performance computing [36].
In many realistic applications, it is either too expensive or intrinsically impossible to collect the full matrix or tensor of measurement or simulation data. That said, the matrix or tensor can be fully reconstructed with reduced numbers of degrees of freedom given the following assumptions: (1) the matrix or tensor exhibits certain algebraic structures such as low-rankness or sparsity, and (2) the observed entries satisfy certain incoherence or randomness. Specifically, the matrix or tensor can be represented as a low-rank matrix or tensor decomposition, such as CANDECOMP/PARAFAC (CP) [32, 44], Tucker [19], tensor train (TT) [70], hierarchical Tucker [31], and tensor ring [92] decompositions. One can then define objective functions as the difference between the observed entries and the extracted ones from the low-rank representation, and use optimization algorithms to update the low-rank factors until convergence. Popular optimization algorithms include nuclear norm minimization [8], alternating least squares (ALS) [3, 44], gradient-based optimization [43], and Riemannian optimization [46], etc. It’s worth mentioning that unlike other low-rank reconstruction problems that allow e.g., randomized sketching of all entries or evaluation of arbitrary entries, the completion problems can only work with a predetermined subset of entries.
@@
For highly oscillatory operators such as large-scale Radon transforms in X-ray tomography or Green’s functions for high-frequency seismic inversion, the aforementioned low-rank matrix and tensor completion algorithms do not approximate the underlying structure of these operators efficiently, leading to increasing rank(s) and high computational cost. Fortunately, butterfly matrices [65, 51, 60, 50, 6] and tensors [42] are particularly designed for highly oscillatory operators. They have been used to represent integral transforms [91, 4, 85, 66, 88], integral operators [65, 30, 58, 75, 59], PDE operators, [57, 12] and neural network parameters [40, 16, 49, 28, 52, 24, 10, 15, 17, 56]. Butterfly decompositions are well-suited to compress measurement operators of sizes $n\times n$ in X-ray and seismic applications with $\mathcal{O}(n\log n)$ (via butterfly matrix) or even $\mathcal{O}(n)$ (via butterfly tensor of [42]) computational and memory complexities. Therefore, butterfly decompositions will be powerful tools if used in matrix or tensor completion of oscillatory operators, as they require significantly lower complexity and less number of observed entries than other existing low-rank completion algorithms. This motivates us to develop completion algorithms in the butterfly format in this current paper.
@@
In this paper, we develop efficient algorithms to perform completion of an oscillatory matrix via butterfly decomposition. We first reshape the matrix into a higher dimensional tensor by using similar tensorization techniques as quantized tensor train (QTT) decomposition [41]. Under such reshaping, the corresponding butterfly matrix decomposition can be converted into a higher-dimensional tensor network, which we call tensor reformulation of butterfly matrices. Note that the result of tensor reformulation is mathematically equivalent to butterfly matrix representation, but makes it convenient to leverage existing tensor network completion software and kernels. This is not to be confused with the butterfly tensor of [42] which directly constructs a tensor decomposition from a tensor operator without reshaping. Next, we propose two tensor completion algorithms for the tensorized butterfly matrix based on ALS and gradient optimization, with an initial guess generated and converted from low-rank matrix completion algorithms. We analyze and demonstrate that for highly oscillatory operators, in the experiments presented in this manuscript, these proposed completion algorithms require only $\mathcal{O}(n\log n)$ observed entries and thus achieve $\mathcal{O}(n\log^{3}n)$ computational complexities. This is in stark contrast to other completion algorithms such as QTT completion or low-rank matrix completion, which require asymptotically larger numbers of entries and arithmetic operations. Finally, we apply the proposed butterfly completion algorithms to several operators in seismic imaging applications to numerically demonstrate their superior convergence, accuracy and efficiency compared with other existing completion algorithms. Aside from applications in tensor completion, we also argue that the proposed tensor reformulation of butterfly can be used to represent neural network layers and allows for efficient automatic differentiation.
@@ 2.1 Tensor Decompositions
Computing an optimal rank tensor decomposition for most problems is NP-hard [34]. Therefore, numerical optimization techniques are used to compute the tensor decomposition by formulating an objective function that is minimized with respect to the tensor decomposition parameters. We formally define the low-rank tensor decomposition optimization problem for a given tensor $\boldsymbol{\mathcal{T}}$ in this subsection. The objective that is minimized to compute the low rank tensor decomposition $\boldsymbol{\mathcal{X}}(\mathcal{S})$,
@@
Various types of tensor decompositions have been proposed such as CANDECOMP/PARAFAC (CP) [32], Tucker [19], tensor train [70], hierarchical Tucker [31], tensor ring [92]. We review the tensor train decomposition in this paper as it has been shown to approximate oscillatory operators [20, 11, 42] and is considered as one of the baseline algorithms in this work. Given an order $2L+2$ tensor $\boldsymbol{\mathcal{T}}$ representing a $L+1$ dimensional operator, the tensor train decomposition or matrix product operator (MPO) representation [35] of $\boldsymbol{\mathcal{T}}$ is written as
@@
Here $i_{0},\ldots,i_{L}$ and $j_{0},\ldots,j_{L}$ are used to index the input and output dimensions of the operator $\boldsymbol{\mathcal{T}}$, and the parameter tensor set is $\mathcal{S}=\{\boldsymbol{\mathcal{S}}^{1},\boldsymbol{\mathcal{S}}^{2},\ldots,\boldsymbol{\mathcal{S}}^{L+1}\}$. Note that the tensors $\boldsymbol{\mathcal{S}}^{1}$ and $\boldsymbol{\mathcal{S}}^{L+1}$ are order $3$ tensors, and the rest are order $4$ tensors. When given an operator in the matrix form, e.g., $\bm{T}\in\mathbb{R}^{n\times n}$ with typically $n=2^{L+1}$, the matrix can be reshaped and permuted into an order $2L+2$ tensor and then represented by the tensor train decomposition in Eq. 3. This is called the quantized tensor train decomposition (QTT) [41] and has been used to represent both sparse and dense matrices [69, 68]. Tensor train decomposition (specifically MPO) has been widely used in accelerating quantum chemistry simulations [39], quantum circuit simulations [27], and many time-dependent, steady-state, low-dimensional and high-dimensional PDEs [26, 13, 63, 89, 22].
@@ 2.2 Tensor Completion
where $\Omega$ is the set of observed entries, and $\mathcal{P}_{\Omega}$ is a selection operator that selects the entries of tensor denoted by $\Omega$. When $\Omega$ is a strict subset of all the entries of the tensor, the computation of the tensor decomposition and using the tensor decomposition to compute the unobserved entries of the tensor is commonly referred to as tensor completion. Low rank matrix completion [37, 33, 86] is one of the most established subjects in this category. Besides low rank matrix completion, many algorithms have been proposed for tensor completion in different formats, such as CP [55], Tucker [25, 46], tensor train [29, 5, 84], hierarchical Tucker [14, 73], and tensor ring decomposition [87].
It is important to note that tensor completion differs from scenarios where arbitrary entries of the tensor can be sampled at will. Instead, only a fixed subset of entries are provided. The arbitrary sampling problem is generally easier than the completion setting, and many algorithms have been proposed for different tensor decompositions in this context as well [67, 74, 21].
@@ 2.3 Software for Tensor Decomposition and Completion
Tensor Toolbox [1] is a MATLAB-based library that provides a wide variety of tensor operations and decompositions for both dense and sparse data. It supports popular models such as CP, Tucker, and block term decompositions. Due to its MATLAB foundation, Tensor Toolbox is widely used for prototyping and algorithm development. However, it is limited to single-node computation and thus is not suitable for very large-scale problems.
Cyclops Tensor Framework (CTF) [83] is a C++ library that provides software infrastructure to perform dense and sparse large-scale tensor computations in a distributed-memory setting. CTF has been used to implement various tensor decomposition algorithms [62, 61, 77, 78]. Recently, new sparse tensor kernels to support tensor completion algorithms in Python were added to CTF [79]. Additionally, a framework to support sparse tensor contractions with tensor networks [38] has been proposed using CTF.
SPLATT (Sparse Tensor Algebra Toolbox) [82] is a high-performance C/C++ library specifically optimized for sparse tensors. It provides parallel implementations of least squares CP decomposition as well as CP completion algorithms and supports modes of parallelism including shared-memory (OpenMP), distributed-memory (MPI), and hybrid configurations. SPLATT uses a compressed sparse fiber (CSF) [80] data structure that generalizes compressed sparse row (CSR) format to tensors.
TuckerMPI is a scalable C++/MPI package designed for large-scale tensor compression based on the Tucker decomposition [2]. It is optimized for distributed computing environments and allows efficient storage and analysis of large-scale dense tensors and grid datasets, making it suitable for large-scale Tucker decomposition.
TensorLy is a Python library that provides high-level tools and efficient implementations for tensor decomposition, tensor algebra, and tensorized deep learning, with extensive backend support for frameworks such as NumPy, PyTorch, and TensorFlow [45]. Many algorithms for CP, Tucker, tensor train decomposition are available in Tensorly with different backends, allowing for efficient and easy-to-use tensor decomposition on different hardware.
@@ 3.1 Complementary Low Rank Property
This discretized matrix may be dense and high-rank, however, it exhibits a structured pattern in which certain submatrices are numerically low-rank. The property that characterizes which blocks are low-rank is known as the complementary low-rank property (CLR) [51]. This property is mathematically described as follows.
@@
where the rank of each submatrix $\bm{T}(\tau^{l}_{i},\nu^{L-l}_{j})$, denoted by $r^{l}_{i,j}$, is bounded by a small constant $r$ (the butterfly rank)[60, 50, 51, 72]. This special property enables the computation a butterfly decomposition, which can reduce the storage complexity of a matrix and computational complexity of matrix-vector multiplications from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$. We describe the butterfly decomposition and briefly describe the algorithms to compute it in the next subsection.
@@ 3.2 Butterfly Matrix Decomposition
When any entry of $\bm{T}$ can be sampled in $\mathcal{O}(1)$, a butterfly decomposition can be efficiently approximated using SVD [51] in $\mathcal{O}(n^{1.5})$, or using Chebyshev interpolation [7, 72] or proxy points [64, 30, 65] in $\mathcal{O}(n\log n)$ complexity. When $\bm{T}$ can be applied to arbitrary random vectors in $\mathcal{O}(n\log n)$ complexity, a butterfly decomposition can be approximated using sketching algorithms [60] in $\mathcal{O}(n^{1.5}\log n)$ complexity. Note that none of these methods can be applied when only a fixed subset of entries is given (i.e., the matrix completion setting). In this setting, numerical optimization algorithms must be used and are described in Section 5.

### v3

@@ Abstract
We address the problem of recovering highly oscillatory operators, represented as $n\times n$ matrices with a fixed set of observed entries. Given that these matrices can be well compressed by butterfly matrix decomposition of $L=\mathcal{O}(\log n)$ levels requiring only $\mathcal{\mathcal{}}{O}(n\log n)$ degrees of freedom, we propose a novel reformulation of the butterfly structure as a compact tensor network. Specifically, we reshape the input matrix as an order $2L+2$ dense tensor, and cast its butterfly decomposition as a tensor network consisting of order $L+3$ dense tensors. This enables efficient utilization of the existing software infrastructure for dense and sparse tensor computations. Next, we propose several tensor completion algorithms based on the tensor reformulation of butterfly format, and compare them against algorithms using the quantized tensor train (QTT) format. These algorithms leverage popular completion methods such as alternating least squares, gradient-based optimization, and alternating directions fitting. Moreover, we propose a novel strategy that uses low-rank matrix completion to efficiently generate an initial guess for the proposed algorithms. To demonstrate the efficiency and applicability of our proposed algorithms, we perform three numerical experiments using simulated oscillatory operators in seismic applications. In these experiments, we use $\mathcal{O}(n\log n)$ observed entries in the input matrix and demonstrate an $\mathcal{O}(n\log^{3}n)$ computational cost of the proposed algorithms, leading to speedups of orders of magnitude and significant accuracy improvements for large matrices compared to state-of-the-art low-rank matrix and QTT tensor completion algorithms.
@@ 1 Introduction
Matrix or tensor completion is the task of inferring missing entries in 2D matrices or higher-order tensors from partial observations, which may be corrupted by noise. Completion problems have been widely studied for applications across data science (e.g., recommendation system [dzhenzher2025low] and computer vision [liu2012tensor]), imaging science (e.g., seismic inversion [9785826, Kumar2015efficient], magnetic resonance imaging [yi2021joint, lingala2011accelerated], and X-ray tomography [candes2013matrixcompletion]), and high performance computing [hutter2022high].
More recent advances involve integrating more sophisticated low-rank structures and smooth data priors into the completion algorithm frameworks
@@
For highly oscillatory operators such as large-scale Radon transforms in X-ray tomography or Green’s functions for high-frequency seismic inversion, the aforementioned low-rank matrix and tensor completion algorithms do not approximate the underlying structure of these operators efficiently, leading to increasing rank(s) and high computational cost. Fortunately, butterfly matrices [michielssen_multilevel_1996, li2015butterfly, liu2021butterfly, Yingzhou_2017_IBF, candemyin09] and tensors [kielstra2024linear] are particularly designed for highly oscillatory operators. They have been used to represent integral transforms [Lexing_SFT_2009, bremer2021rapid, Tygert_2010_spherical, Oneil_2010_specialfunction, Haizhao_2018_Phase], integral operators [michielssen_multilevel_1996, Han_2017_butterflyLUPEC, Liu_2017_HODBF, Sadeed2022VIE, liu2023fast], PDE operators, [liu2021sparse, claus2023sparse] and neural network parameters [Khoo2020switchnet, dao2019learning, li2022wide, gonon2024make, DEBUT, fan2022adaptablebutterflyacceleratorattentionbased, chen2022pixelated, pmlr-v162-dao22a, dao2020kaleidoscope, liu2023parameter]. Butterfly decompositions are well-suited to compress measurement operators of sizes $n\times n$ in X-ray and seismic applications with $\mathcal{O}(n\log n)$ (via butterfly matrix) or even $\mathcal{O}(n)$ (via butterfly tensor of [kielstra2024linear]) computational and memory complexities. Therefore, butterfly decompositions will be powerful tools if used in matrix or tensor completion of oscillatory operators, as they require significantly lower complexity and less number of observed entries than other existing low-rank completion algorithms. This motivates us to develop completion algorithms in the butterfly format in this current paper.
@@
In this paper, we develop efficient algorithms to perform completion of an oscillatory matrix via butterfly decomposition. Our contribution is three-fold: (1) First, we reshape the matrix into a higher dimensional tensor by using similar tensorization techniques as quantized tensor train (QTT) decomposition [khoromskij2011d]. Under such reshaping, the corresponding butterfly matrix decomposition can be converted into a higher-dimensional tensor network, which we call tensor reformulation of butterfly matrices. Note that the result of tensor reformulation is mathematically equivalent to butterfly matrix representation, but makes it convenient to leverage existing tensor network completion software and kernels. This is not to be confused with the butterfly tensor of [kielstra2024linear] which directly constructs a tensor decomposition from a tensor operator without reshaping. (2) Next, we propose three tensor completion algorithms for the tensorized butterfly matrix based on ALS, ADF and gradient optimization, with an initial guess generated and converted from low-rank matrix completion algorithms. We analyze and demonstrate that for highly oscillatory operators, in the experiments presented in this manuscript, these proposed completion algorithms require only $\mathcal{O}(n\log n)$ observed entries and thus achieve $\mathcal{O}(n\log^{3}n)$ computational complexities. This is in stark contrast to other completion algorithms such as QTT completion or low-rank matrix completion, which require asymptotically larger numbers of arithmetic operations. It is also worth mentioning that although ALS and gradient optimization-based algorithms have been considered for butterfly matrices [le2025butterflyfactorizationerrorguarantees, DEBUT, dao2019learning] , these existing algorithms assume $n^{2}$ observed entries instead of a matrix completion setting. (3) Finally, we apply the proposed butterfly completion algorithms to several operators in seismic imaging applications to numerically demonstrate their superior convergence, accuracy and efficiency compared with other existing completion algorithms. Aside from applications in tensor completion, we also argue that the proposed tensor reformulation of butterfly can be used to represent neural network layers and allows for efficient automatic differentiation.
@@ 2.1 Tensor Decompositions
Computing an optimal rank tensor decomposition for most problems is NP-hard [hillar2013most]. Therefore, numerical optimization techniques are used to compute the tensor decomposition by formulating an objective function that is minimized with respect to the tensor decomposition parameters. We formally define the low-rank tensor decomposition optimization problem for a given tensor $\boldsymbol{\mathcal{T}}$ in this subsection. The objective that is minimized to compute the low rank tensor decomposition $\boldsymbol{\mathcal{X}}(\mathcal{S})$,
@@
Various types of tensor decompositions have been proposed such as CANDECOMP/PARAFAC (CP) [harshman1970foundations], Tucker [de2000multilinear], tensor train [oseledets2011tensor], hierarchical Tucker [hackbusch2009new], tensor ring [zhao2016tensor]. We review the tensor train decomposition in this paper as it has been shown to approximate oscillatory operators [dolgov2012superfast, chen2024direct, kielstra2024linear] and is considered as one of the baseline algorithms in this work. Given an order $2L+2$ tensor $\boldsymbol{\mathcal{T}}$ representing a $L+1$ dimensional operator, the tensor train decomposition or matrix product operator (MPO) representation [hubig2017generic] of $\boldsymbol{\mathcal{T}}$ is written as
@@
Here $i_{0},\ldots,i_{L}$ and $j_{0},\ldots,j_{L}$ are used to index the input and output dimensions of the operator $\boldsymbol{\mathcal{T}}$, and the parameter tensor set is $\mathcal{S}=\{\boldsymbol{\mathcal{S}}^{1},\boldsymbol{\mathcal{S}}^{2},\ldots,\boldsymbol{\mathcal{S}}^{L+1}\}$. Note that the tensors $\boldsymbol{\mathcal{S}}^{1}$ and $\boldsymbol{\mathcal{S}}^{L+1}$ are order $3$ tensors, and the rest are order $4$ tensors. When given an operator in the matrix form, e.g., $\bm{T}\in\mathbb{R}^{n\times n}$ with typically $n=2^{L+1}$, the matrix can be reshaped and permuted into an order $2L+2$ tensor and then represented by the tensor train decomposition in Eq. 3. This is called the quantized tensor train decomposition (QTT) [khoromskij2011d] and has been used to represent both sparse and dense matrices [oseledetsmatrixapproximation, oseledetstensorinv]. Tensor train decomposition (specifically MPO) has been widely used in accelerating quantum chemistry simulations [keller2015efficient], quantum circuit simulations [gelss2022low], and many time-dependent, steady-state, low-dimensional and high-dimensional PDEs [gelss2022solving, CORONA2017145, manzini2023tensor, ye2024quantized, dolgov2012fast].
@@ 2.2 Tensor Completion
where $\Omega$ is the set of observed entries, and $\mathcal{P}_{\Omega}$ is a selection operator that selects the entries of tensor denoted by $\Omega$. When $\Omega$ is a strict subset of all the entries of the tensor, the computation of the tensor decomposition and using the tensor decomposition to compute the unobserved entries of the tensor is commonly referred to as tensor completion. Low rank matrix completion [jain2013low, hastie2015matrix, vandereycken2013low] is one of the most established subjects in this category. Besides low rank matrix completion, many algorithms have been proposed for tensor completion in different formats, such as CP [liu2012tensor], Tucker [filipovic2015tucker, kressner2014low], tensor train [grasedyck2015variants, cai2022provable, steinlechner2016riemannian], hierarchical Tucker [da2015optimization, rauhut2015tensor], and tensor ring decomposition [wang2017efficient].
It is important to note that tensor completion differs from scenarios where arbitrary entries of the tensor can be sampled at will. Instead, only a fixed subset of entries are provided. The arbitrary sampling problem is generally easier than the completion setting, and many algorithms have been proposed for different tensor decompositions in this context as well [oseledets2010tt, saibaba2016hoid, dolgov2021functional].
@@ 2.3 Software for Tensor Decomposition and Completion
Tensor Toolbox [bader2008efficient] is a MATLAB-based library that provides a wide variety of tensor operations and decompositions for both dense and sparse data. It supports popular models such as CP, Tucker, and block term decompositions. Due to its MATLAB foundation, Tensor Toolbox is widely used for prototyping and algorithm development. However, it is limited to single-node computation and thus is not suitable for very large-scale problems.
Cyclops Tensor Framework (CTF) [ctf] is a C++ library that provides software infrastructure to perform dense and sparse large-scale tensor computations in a distributed-memory setting. CTF has been used to implement various tensor decomposition algorithms [ma2018accelerating, ma2021efficient, singh2021comparison, singh2023alternating]. Recently, new sparse tensor kernels to support tensor completion algorithms in Python were added to CTF [singh2022distributed]. Additionally, a framework to support sparse tensor contractions with tensor networks [kanakagiri2024minimum] has been proposed using CTF.
SPLATT (Sparse Tensor Algebra Toolbox) [smith2015splatt] is a high-performance C/C++ library specifically optimized for sparse tensors. It provides parallel implementations of least squares CP decomposition as well as CP completion algorithms and supports modes of parallelism including shared-memory (OpenMP), distributed-memory (MPI), and hybrid configurations. SPLATT uses a compressed sparse fiber (CSF) [smith2015tensor] data structure that generalizes compressed sparse row (CSR) format to tensors.
TuckerMPI is a scalable C++/MPI package designed for large-scale tensor compression based on the Tucker decomposition [Ballard2019TuckerMPI]. It is optimized for distributed computing environments and allows efficient storage and analysis of large-scale dense tensors and grid datasets, making it suitable for large-scale Tucker decomposition.
TensorLy is a Python library that provides high-level tools and efficient implementations for tensor decomposition, tensor algebra, and tensorized deep learning, with extensive backend support for frameworks such as NumPy, PyTorch, and TensorFlow [Kossaifi2019TensorLy]. Many algorithms for CP, Tucker, tensor train decomposition are available in Tensorly with different backends, allowing for efficient and easy-to-use tensor decomposition on different hardware.
@@ 3.1 Complementary Low Rank Property
This discretized matrix may be dense and high-rank, however, it exhibits a structured pattern in which certain submatrices are numerically low-rank. The property that characterizes which blocks are low-rank is known as the complementary low-rank property (CLR) [li2015butterfly]. This property is mathematically described as follows.
@@
where the rank of each submatrix $\bm{T}(\tau^{l}_{i},\nu^{L-l}_{j})$, denoted by $r^{l}_{i,j}$, is bounded by a small constant $r$ (the butterfly rank)[liu2021butterfly, Yingzhou_2017_IBF, li2015butterfly, Pang2020IDBF]. This special property enables the computation a butterfly decomposition, which can reduce the storage complexity of a matrix and computational complexity of matrix-vector multiplications from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$. We describe the butterfly decomposition and briefly describe the algorithms to compute it in the next subsection.
@@ 3.2 Butterfly Matrix Decomposition
When any entry of $\bm{T}$ can be sampled in $\mathcal{O}(1)$, a butterfly decomposition can be efficiently approximated using SVD [li2015butterfly] in $\mathcal{O}(n^{1.5})$, or using Chebyshev interpolation [candes2007fast, Pang2020IDBF] or proxy points [Eric_1994_butterfly, Han_2017_butterflyLUPEC, michielssen_multilevel_1996] in $\mathcal{O}(n\log n)$ complexity. When $\bm{T}$ can be applied to arbitrary random vectors in $\mathcal{O}(n\log n)$ complexity, a butterfly decomposition can be approximated using sketching algorithms [liu2021butterfly] in $\mathcal{O}(n^{1.5}\log n)$ complexity. Note that none of these methods can be applied when only a fixed subset of entries is given (i.e., the matrix completion setting). In this setting, numerical optimization algorithms must be used and are described in Section 5.